OpenStudy (goformit100):

Admission to the zoo costs $4 for adults and $2 for kids. The total bill for your entire extended family was $30. This scenario can be modeled by the equation 4a + 2k = 30. Which is not a possible combination of adults and kids? 5 adults and 5 kids 4 adults and 7 kids 7 adults and 1 kid 3 adults and 6 kids

OpenStudy (goformit100):

@shubhamsrg

OpenStudy (ivettef365):

just substitute the numbers on the equation to see which one is not true

OpenStudy (goformit100):

How ? Give me the perfect answer

OpenStudy (ivettef365):

for example you have equation 4a + 2k = 30 5 adults and 5 kids 4(5) + 2(5) = 30 20 + 10 = 30 30 = 30 true so that is ok

OpenStudy (ivettef365):

so find the one that is NOT possible

OpenStudy (goformit100):

So, what will be the answer ?

OpenStudy (ivettef365):

ok, since you have 99 I guess I can give you the answer 3 adults and 6 kids 4(3) + 2(6) = 30 12 + 12 = 30 24 = 30 NOT GOOD

OpenStudy (goformit100):

Thankyou sir

OpenStudy (ivettef365):

madam